OpenStudy (anonymous):

Simplify x^2 - 9 / 2x^2 - 3x -9

OpenStudy (anonymous):

Sorry I'm trying to write this out in proper equation

OpenStudy (saifoo.khan):

\[\frac{ x^2 - 9}{2x^2 - 3x -9}\]

OpenStudy (saifoo.khan):

Like that right?

OpenStudy (anonymous):

Yes that's it

OpenStudy (anonymous):

Couldn't dot he divide thing

OpenStudy (saifoo.khan):

\[\frac{(x-3)(x+3)}{(x-3) (2 x+3)}\] \[\frac{\cancel{(x-3)}(x+3)}{\cancel{(x-3)} (2 x+3)}\]

OpenStudy (saifoo.khan):

\[\frac{x+3}{2x+3}\]

OpenStudy (anonymous):

Could you explain like step by step, sorry I'm not very good at this

OpenStudy (saifoo.khan):

What u have to do is, u have to factor the both numerator and denominator..

OpenStudy (saifoo.khan):

After factoring, u will get, \[\frac{(x-3)(x+3)}{(x-3) (2 x+3)}\] Now u have to cancel the like terms, leaving u with, \[\frac{x+3}{2x+3}\]

OpenStudy (anonymous):

Thankyou I get it now, as soon as you factorised it's easy from there thanks a lot!

OpenStudy (saifoo.khan):

No problem! ^_^